DFB, Inc. expects earnings next year of $5.01 per share, and it plans to pay a $3.42 dividend to shareholders (assume that is one year from now). DFB will retain $1.59 per share of its earnings to reinvest in new projects that have an expected return of 15.5% per year. Suppose DFB wil maintain the same idend payout rate, retention rate, and return on new investments in the future and will not change its number of outstanding shares. Assume next dividend is due in one year. a. What growth rate of earnings would you forecast for DFB? b.H DFBs equity cost of capital is 11.1%, what price would you estimate for DFB stock? c. Suppose instead that DFB paid a dividend of $4.42 per share at the end of this year and retained only $0.59 per share in earnings. That is, it chose to pay a higher dividend instead of reinvesting in as many new projects. If DFB maintains this higher payout rate in the future, what stock price would you estimate for the firm now? Should DFB raise its dividend? a. What growth rate of earnings would you forecast for DFB? DFBs growth rate of earnings is LM. (Round to one decimal place.)
Answer #1

DFB Inc:

Earnings per share next year, E1 = $5.01

Dividend per share, next year D1 = $3.42

Retention ratio, b = $1.59/$5.01 = 0.32

Payout ratio (1-b) = DPS/EPS = 3.42/5.01 = 0.68

Expected return on equity, ROE= 15.5% per year

a) growth rate of earnings forecast, g = ROE * b = 0.155 * 0.32 = 0.0496 = 4.96%

b)If cost of equity Ke = 11.10%, then Price = D1 / (Ke - g) = $3.42 / (0.111 - 0.0496) = $3.42/0.0614 = $55.70

c)Paid dividend of $4.42, then retention ratio, b = (1 - DPS/EPS) = ( 1 - $4.42/$5.01) = (1 - 0.88) = 0.12

  • g = ROE * b = 0.155 * 0.12 = 0.0186 = 1.86%
  • Stock price, P = D1 / (Ke - g)
  • P = $4.42 / (0.111 - 0.0186)
  • P = $4.42 / 0.0924
  • P = $47.84

Usually, companies increase dividends when there are no positive NPV projects available. Maintaining increased dividends is a costly affair, as, decrease or cut of dividends in future signals negative growth prospects about the company earnings potential.

Company is better off in investing in new positive NPV projects, as it helps grow business at higher rate of returns, thus a potential upside in the stock price. When dividend was increased, the stock price reduced to $47.84 from $55.70.

So it is better to reinvest than not to increase the dividend.